S=180n-360
solve for N

Answer by mananth(16949)

S=180n-360
add 360 to both sides
s+360 = 180n-360+360
s+360=180n
divide by 180
(s+360)/180 = 180n/180
s/180 + 360 / 180 = n
s/180 + 2 = n